# PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks on Non-IID Data

### v4

@@ Abstract
Privacy-Preserving Federated Learning (PPFL) allows multiple clients to collaboratively train by submitting hidden model updates. Nonetheless, PPFL is vulnerable to data poisoning attacks due to the distributed training nature of clients in cross-silo scenarios. Existing solutions have struggled to improve the performance of PPFL in poisoned Non-IID data. To address the issues, this paper proposes a privacy-preserving federated prototype learning framework, named PPFPL, which enhances the cross-silo FL performance in poisoned Non-IID data while protecting the privacy of clients. Specifically, we adopt prototypes as client-submitted model updates to eliminate the impact of poisoned data distribution. In addition, we design a secure aggregation protocol by utilizing homomorphic encryption to achieve Byzantine-robust aggregation on two servers, which greatly reduces the impact of malicious clients. Theoretical analyses confirm the convergence and privacy of PPFPL, and experimental results on publicly datasets show that PPFPL is effective for resisting data poisoning attacks with Non-IID data.
Index Terms:
@@ I Introduction
Federated Learning (FL) is a distributed learning paradigm where each client shares its model updates instead of raw training data. In industrial applications, massive data is distributed across independent organizations governed by strict privacy regulations [1]. To break data silos among organizations without compromising privacy, cross-silo FL provides a viable solution for industrial scenarios [2][3]. Specifically, in cross-silo FL, the clients are usually large organizations, resulting in a relatively small number of clients with significant computational capabilities. However, client-submitted model updates are vulnerable to privacy attacks, which threats the privacy security of cross-silo FL [4][5]. To mitigate the privacy risks, Differential Privacy (DP)-based [6][7] and Homomorphic Encryption (HE)-based [8] [9] PPFL approaches are proposed. In particular, DP-based approach is commonly applied to cross-device FL due to its low computation overhead, but degrades the FL performance by introducing noise. Conversely, HE-based approach offers higher privacy security without sacrificing the global model accuracy, thus making it more suitable for cross-silo FL.
@@
While HE-based approaches demonstrated their effectiveness in terms of privacy preservation, they are susceptible to data poisoning attacks [10][11]. Specifically, malicious clients launch data poisoning attacks by tampering with their raw training data, and submit model updates derived from the poisoned data, thereby degrading the performance of PPFL [12][13]. Moreover, Moreover, privacy-preserving techniques obscure the model updates from malicious clients, making data poisoning attacks more concealed and more difficult to defend. To audit obscured model updates, existing methods [14, 15, 16, 17] utilize HE and Secure Multi-party Computing (SMC) to identify malicious model updates in ciphertext. However, these works overlook the data heterogeneity in FL, i.e., the data among clients is typically non-independent and identically distributed (Non-IID). The Non-IID data leads to the inconsistency among benign model updates, making it difficult for these defense methods [14, 15, 16, 17] to distinguish whether the deviations are caused by data heterogeneity or by malicious model updates. To distinguish between benign and malicious updates in the Non-IID data, existing methods [18, 19, 20] employ clustering or adaptive aggregation weighting operation to mitigate the inconsistency of model updates during the auditing phase, thereby filtering out updates caused by model poisoning attacks. However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, thereby forming poisoned Non-IID data. These manipulations distort the optimization direction of local training, causing these methods [18, 19, 20] to optimize toward the tampered Non-IID distribution, thereby degrading the performance of FL. Therefore, a critical challenge remains: how to enhance the performance of PPFL in poisoned Non-IID data while effectively resisting data poisoning attacks.
@@
Inspired by prototype learning [21], several works [22, 23, 24, 25] proposed prototype aggregation to address the Non-IID issue in FL, where prototypes are exchanged between the server and clients. Each prototype represents a class-level feature, computed as the mean of the feature representations of samples belonging to the same class. For instance, the work in [23] suggest that in recognizing the “cat”, different clients have their own unique “imaginary picture” or “prototype” to represent the concept of “cat”. By exchanging these prototypes, clients gain more knowledge about the concept of “cat”. Consequently, prototypes, as feature representations independent of data distribution, inspire our work.
@@
Our main contributions are summarized as follows.
  • We are the first to introduce prototype learning into PPFL to defend against data poisoning attacks. By transferring prototypes between clients and servers, our framework mitigates the impact of tampered data distribution and enhances the FL performance in poisoned Non-IID data.
  • We utilize HE and SMC techniques to design a secure aggregation protocol across two servers to filter prototypes submitted by malicious clients, ensuring Byzantine-robust results while protecting the privacy of benign clients.
  • We provide a convergence guarantee for PPFPL in the presence of data poisoning attacks, which theoretically ensures the framework’s feasibility.
The rest of the paper is organized as follows. Section II reviews PPFL works against poisoning attacks. Section III introduces prototype learning and HE. Section IV formalizes the system model. Our PPFPL is detailed in Section V. Section VI provide theoretical analysis. Section VII reviews experimental results. Finally, Section VIII concludes this paper.
II RELATED WORK
Privacy preservation in FL: Although FL has naturally certain privacy protection, it remains vulnerable to privacy attacks, which causes the privacy threat of benign clients. To resist such attacks, DP-based and HE-based approaches are proposed to preserve client-submitted model updates. Specifically, DP-based schemes [26, 27, 28, 29] deploy local differential privacy into model updates, ensuring privacy without compromising the utility of model updates. Despite their low computation overheads, these schemes add Gaussian noise or Laplace noise into local model training, which degrades the FL performance to some extent. In contrast, HE is a commonly used cryptographic primitive in across-silo PPFL that provides strong privacy preservation without sacrificing the global model’s accuracy. Specifically, Fang and Qian are one of the first scholars to implement PPFL using HE [30]. They proposed a multi-party machine learning scheme using Paillier [31] technique without compromising clients’ privacy. Considering the heavy communication overhead of Paillier, the work in [32] proposed a privacy-preserving FL using CKKS (i.e., Cheon-Kim-Kim-Song) that reduces computational overhead associated with ciphertexts. This is because CKKS is more efficient and better suitable to handle large-scale vector and multi-parameter network models than Paillier [33]. However, the aforementioned schemes overlook the threat of data poisoning attacks caused by distributed training.
Resisting data poisoning attacks in FL: Tolpegin et al. demonstrated that data poisoning attacks can significantly reduce the classification accuracy of the global model, even with a small percentage of malicious clients [34]. Additionally, they proposed an aggregated defense strategy that identifies malicious clients in FL to circumvent data poisoning attacks. Similarly, the works in [35, 36, 37] proposed detection methods to filter malicious model updates. Differently, Doku et al. employed SVM to audit client’s local training data for excluding malicious clients [38]. However, this work violates the privacy of clients to some extent. Furthermore, the threats of privacy and data poisoning attacks usually coexist in practice across-silo FL. PPFL approaches can utilize cryptographic primitives to provide privacy preservation for clients, while concealing data poisoning attacks from malicious clients.
Resisting data poisoning attacks in PPFL: Considering the threats of both privacy and data poisoning attacks, the works such as [15, 14, 16, 17] combine HE and SMC to identify anomalous model updates in ciphertexts. For instance, the work in [15] proposed a privacy-enhanced FL framework that adopts HE as the underlying technology and provides two servers with a channel to punish malicious clients via gradient extraction of logarithmic function. Similarly, the work in [14] designed a validity checking protocol for ciphertexts under two servers to protect data privacy and adaptively adjust the weight of clients’ gradients to weaken data poisoning attacks. However, these schemes [15, 14, 16, 17] ignore a fundamental problem, i.e., they neglect the deviations from model updates caused by Non-IID data, making it difficult to distinguish whether the deviations come from malicious model updates or Non-IID data.
To audit malicious updates in PPFL with Non-IID data, the works in [18][19] are proposed to eliminate the deviations caused by Non-IID data during the aggregation stage. Specifically, ShieldFL designs a Byzantine-tolerant aggregation mechanism to prevent misjudgments on outliers caused by Non-IID data [18]. Furthermore, Chen et al. adopts clustering combined with cosine similarity and median strategies to eliminate deviations among model updates during aggregation auditing [19]. These schemes can only resist model poisoning attacks confronted by federated learning with Non-IID data, but they cannot essentially improve the performance of FL on Non-IID data. They need to be combined with specialized training techniques designed for Non-IID data (e.g., FedProx[39], FedDyn[40], or FedLC[41], etc.) to radically improve performance of FL. These specialized techniques adds an auxiliary term into local optimization function to constrain model updates of clients, which helps to increase consistency of model updates during local model training. However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, creating poisoned Non-IID data, which causes specialized training techniques to constrain model updates based on these compromised inputs. Therefore, the above works [18][19] cannot be combined with the specialized techniques to resist data poisoning attacks while enhancing the FL performance under poisoned Non-IID data.
Federated prototype learning: Recently, prototype learning is gradually being applied in federated learning to solve the Non-IID issue. Specifically, the works in [22, 23, 24, 25] are one of the first to propose federated prototype learning using the concept of prototype learning. Different from the specialized training techniques (e.g., FedProx, FedDyn, or FedLC), the core idea enables clients to pull the same-class samples towards the global prototypes of that class and away from the global prototypes of other classes. In other words, each class holds its corresponding prototype that is independent of other classes. As a result, client-submitted prototype is affected by the samples but is independent of data distribution among clients. This inspires our work: federated prototype learning allows distributed training to remain independent of tampered data distributions caused by data poisoning attacks.
Although prototype learning has already been studied to improve the performance of federated learning in Non-IID data, it has not been explored in privacy-preserving federated learning against data poisoning attacks. Our work is the first to integrate prototype learning into PPFL to tackle the performance degradation caused by tampered Non-IID data in across-silo FL, while preserving the privacy of client-submitted prototypes.
III PRELIMINARIES
This section introduces prototypes in federated learning and CKKS technology.
III-A Prototypes Meet Federated Learning
In the classification task of prototype learning, the prototype is defined as a feature vector representing a specific class [42]. Due to this property, the prototypes of the same class among clients are similar in FL task. Consequently, many FL schemes [22, 23, 24, 25] enhance the handling of Non-IID data by prototype learning, which allows clients to align their prototypes with other clients during local model training.
To further understand the prototype calculation in federated prototype learning, we introduce some basic notations below. Let $\mathcal{S}$ be the set of clients, where each client $m\in\mathcal{S}$ has an independent private dataset, denoted as $\mathcal{D}_{m}=\{(\textit{{x}}_{(i)},y_{(i)})\}^{|\mathcal{D}_{m}|}$. Here, $|\mathcal{D}_{m}|$ represents the number of samples in client $m$, and $(\textit{{x}}_{(i)},y_{(i)})$ denotes sample $i$ in dataset, where $\textit{{x}}_{(i)}$ and $y_{(i)}$ correspond to the feature vector and class label of sample $i$, respectively. Meanwhile, let $\mathcal{I}$ be the set of classes in classification task, where each class $k$ belongs to $\mathcal{I}$. In classification task, the local model includes a feature extractor and a decision classifier. Specifically, the feature extractor transforms sample features into compressed features, while decision classifier maps the compressed features to get classification results. Formally, let $f_{m}(\textit{{r}}_{m,t};\cdot)$ be feature extractor for client $m$, parameterized by $\textit{{r}}_{m,t}$, where $t$ denotes the $t$-th communication round. Given the feature $\textit{{x}}_{(i)}$ of sample $i$, it is input to feature extractor to obtain compressed feature $\textit{{u}}_{(i)}=f_{m}(\textit{{r}}_{m,t};\textit{{x}}_{(i)})$. Let $g_{m}(\textit{{z}}_{m,t};\cdot)$ be decision classifier for client $m$, parameterized by $\textit{{z}}_{m,t}$. The classifier maps the compressed feature $\textit{{u}}_{(i)}$ to predict the class $y^{\prime}=g_{m}(\textit{{z}}_{m,t};\textit{{u}}_{(i)})$. Thus, we denote the local model as $\mathcal{F}_{m}((\textit{{r}}_{m,t},\textit{{z}}_{m,t});\cdot)=g_{m}(\textit{{z}}_{m,t};\cdot)\circ f_{m}(\textit{{r}}_{m,t};\cdot)$, where $\circ$ denotes composite operator. To simplify notation, we use $\textit{{w}}_{m,t}$ to denote $(\textit{{r}}_{m,t},\textit{{z}}_{m,t})$, so we have $\mathcal{F}_{m}((\textit{{r}}_{m,t},\textit{{z}}_{m,t});\cdot)=\mathcal{F}_{m}(\textit{{w}}_{m,t};\cdot)$, and $\textit{{w}}_{m,t}$ is consider as model parameters for client $m$. Next, we introduce calculation process of prototypes.
In federated prototype learning, prototypes can be categorized into local prototypes, computed by clients, and global prototypes, aggregated by the server. Specifically, each client’s goal is to align its local prototype to global prototype during local model training. Thus, each client computes its local prototype via its training dataset during local model training. Formally, let $\textit{{c}}^{k}_{m,t}$ be the local prototype of class $k\in\mathcal{I}$ at client $m$ in $t$-th communication round, calculated as
$\textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}f_{m}(\textit{{r}}_{m,t};\textit{{x}}_{(i)}),\forall k\in\mathcal{I},$ (1)
where $\mathcal{D}^{k}_{m}$ denotes the dataset with class $k$ at client $m$. The $\textit{{c}}^{k}_{m,t}$ can be understood as the mean of compressed features of samples belonging to class $k$ at client $m$. Further, we can get local prototypes for all classes at client $m$, denoted as $\{\textit{{c}}^{k}_{m,t}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$. Once local model training is complete, each client submits its local prototypes to server.
To calculate global prototype, the server adopts an averaging operation to local prototypes submitted by clients. Thus, the global prototype $\textit{{C}}^{k}_{t+1}$ for each class is calculated as follows:
$\textit{{C}}^{k}_{t+1}$ $=\frac{1}{|\mathcal{S}|}\sum_{m\in\mathcal{S}}\textit{{c}}_{m,t}^{k},\forall k\in\mathcal{I},$ (2)
where $|\mathcal{S}|$ denotes the number of clients. For global prototypes of all classes, we denote them by the set $\{\textit{{C}}^{k}_{t+1}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$. Subsequently, the server distributes the latest global prototypes to each client to further train their local model.
III-B CKKS
CKKS is a homomorphic encryption technology. Similar to traditional HE, it is a cryptographic technique based on mathematical computations, allowing operations on encrypted data without decrypting. However, traditional HE schemes such as RSA [43], ElGamal [44], and Paillier are limited in that they only support either additive or multiplicative operations but not both simultaneously. In contrast, the CKKS technology provides both additive and multiplicative homomorphic encryption, which is known as full HE [45]. In addition, CKKS is known for its efficiency, especially in terms of encryption/decryption speed when dealing with scale vectors of varying parameter lengths. Therefore, we employ CKKS for the privacy protection of clients due to its computational efficiency. For a more detailed implementation principle and process of CKKS, please refer to paper [46].
IV Problem Statement

### v5

@@ Abstract
Privacy-Preserving Federated Learning (PPFL) enables multiple clients to collaboratively train models by submitting secreted model updates. Nonetheless, PPFL is vulnerable to data poisoning attacks due to its distributed training paradigm in cross-silo scenarios. Existing solutions have struggled to improve the performance of PPFL under poisoned Non-Independent and Identically Distributed (Non-IID) data. To address the issues, this paper proposes a privacy-preserving federated prototype learning framework, named PPFPL, which enhances the cross-silo FL performance against poisoned Non-IID data while protecting client privacy. Specifically, we adopt prototypes as client-submitted model updates to eliminate the impact of poisoned data distributions. In addition, we design a secure aggregation protocol utilizing homomorphic encryption to achieve Byzantine-robust aggregation on two servers, significantly reducing the impact of malicious clients. Theoretical analyses confirm the convergence and privacy of PPFPL. Experimental results on public datasets show that PPFPL effectively resists data poisoning attacks under Non-IID settings.
{IEEEkeywords}
@@ 1 Introduction
Federated Learning (FL) is a distributed learning paradigm where each client shares its model updates instead of raw training data. In industrial applications, large volumes of data are distributed across independent organizations governed by strict privacy regulations [1]. To break data silos among organizations without compromising privacy, cross-silo FL provides a viable solution for industrial scenarios [2][3]. Specifically, in cross-silo FL, clients are usually large organizations, resulting in a relatively small number of clients with significant computational capabilities. However, client-submitted model updates are vulnerable to privacy attacks, threatening the security of cross-silo FL [4][5]. To mitigate the privacy risks, Differential Privacy (DP)-based [6][7] and Homomorphic Encryption (HE)-based [8][9] PPFL approaches are proposed. In particular, DP-based approach is commonly applied to cross-device FL due to its low computation overhead, but degrades the FL performance by injecting noise. Conversely, HE-based approach provides higher privacy security without sacrificing the global model accuracy, making it more suitable for cross-silo FL.
@@
While HE-based approaches have demonstrated their effectiveness in terms of privacy preservation, they are susceptible to data poisoning attacks [10][11]. Specifically, malicious clients launch data poisoning attacks by tampering with their raw training data and submitting model updates derived from the poisoned data, thereby degrading the performance of PPFL [12][13]. Moreover, privacy-preserving techniques obscure the model updates from malicious clients, making data poisoning attacks more concealed and difficult to defend. To audit secreted model updates, existing methods [14, 15, 16, 17, 18] utilize HE and Secure Multi-party Computing (SMC) to identify malicious model updates within ciphertext. However, these works overlook the data heterogeneity in FL, i.e., the data among clients is typically non-independent and identically distributed (Non-IID). The Non-IID data leads to the inconsistency among benign model updates, making it difficult for these defense methods [14, 15, 16, 17, 18] to distinguish whether the deviations are caused by data heterogeneity or by malicious manipulation. To distinguish between benign and malicious updates under Non-IID data, existing methods [19, 20, 21] employ clustering or adaptive aggregation weighting operation to mitigate the inconsistency of model updates during the auditing phase. However, in data poisoning attacks, malicious clients manipulate the features and labels of their training data, thereby forming poisoned Non-IID data. These manipulations distort the optimization direction of local training, causing these methods [19, 20, 21] to optimize toward the tampered Non-IID distribution, severely degrading the performance of FL. Therefore, a critical challenge remains: how to enhance the performance of PPFL under poisoned Non-IID data while effectively resisting data poisoning attacks.
@@
Inspired by prototype learning [22], several works [23, 24, 25, 26] have introduced prototype aggregation to address the Non-IID problem in FL by exchanging prototypes between servers and clients. Each prototype represents a class-level feature, computed as the mean of the feature representations of samples within the same class. For instance, the work in [25] suggests that in recognizing the class “cat,” different clients have their own unique “imaginary picture” or “prototype” to represent the concept of “cat”. By prototype exchange, clients gain more knowledge about the concept of “cat”. Consequently, prototypes, as feature representations independent of local data distributions, motivate our work.
In cases of poisoned Non-IID data, both tampered data features and distributions cannot be repaired since the server cannot control malicious client behavior. This prompts us to pose a question: Is it possible to design a PPFL that leverages prototype learning, so that client-submitted model updates are affected only by tampered data features rather than by tampered data distribution, while incorporating a secure aggregation protocol to eliminate the impact of tampered data features, thereby achieving Byzantine-robust results (i.e., guaranteeing reliability under arbitrary and potentially malicious behaviors in distributed computing systems [27])? To answer this question, we propose a Privacy-Preserving Federated Prototype Learning framework, named PPFPL, which is suitable for cross-silo scenarios. The PPFPL framework consists of two “non-colluding” servers (assumed to be honest but curious) and multiple clients. The key novelty of PPFPL lies in: $(\textbf{\text{i}})$ We propose a novel local optimization function leveraging prototype learning, where each client submit prototypes to two servers, which mitigates the impact of tampered data distributions. Unlike existing prototype-based works [23][24], our optimization function employs cosine similarity to constrain the influence of malicious clients (see the proof of Theorem 1). $(\textbf{\text{ii}})$ We design a secure aggregation protocol between two servers to aggregate client-submitted prototypes using HE and SMC techniques. This protocol achieves Byzantine-robust results while preserving the privacy of benign clients.
Our main contributions are summarized as follows.
  • To the best of our knowledge, this is the first work that introduces prototype learning into PPFL to defend against data poisoning attacks. By transferring prototypes between clients and servers, our framework mitigates the impact of tampered data distribution and enhances the FL performance in poisoned Non-IID data.
  • We employ HE and SMC techniques to design a secure aggregation protocol across two servers that filters malicious prototypes, thereby ensuring Byzantine-robust results while protecting the privacy of benign clients. Notably, existing studies have not provided solutions for privacy protection in prototype-based federated learning, and our work fills this important research gap.
  • We establish theoretical guarantees on convergence and privacy for PPFPL, thereby ensuring the framework’s feasibility.
The rest of the paper is organized as follows. Section 2 reviews PPFL works against poisoning attacks. Section 3 introduces prototype learning and HE. Section 4 formalizes both the system model and threat model. Our PPFPL is detailed in Section 5. Section 6 provides theoretical analysis. Section 7 reviews experimental results. Finally, Section 8 concludes this paper.
2 RELATED WORK
Privacy preservation in FL: Although FL inherently provides a degree of privacy protection, it remains vulnerable to privacy attacks, which causes the privacy threat of benign clients. To resist such attacks, DP-based and HE-based approaches have been proposed to preserve client-submitted model updates. Specifically, DP-based schemes [28, 29, 30, 31] deploy local differential privacy into model updates, ensuring privacy without compromising the utility of model updates. Despite their low computation overheads, these schemes introduce Gaussian noise or Laplace noise into local model training, which inevitably degrades the FL performance to some extent. In contrast, HE is a commonly adopted cryptographic primitive in across-silo PPFL that provides strong privacy preservation without sacrificing the global model’s accuracy. Fang and Qian are one of the first scholars to implement PPFL using HE [32]. They proposed a multi-party machine learning scheme using Paillier [33] technique without compromising clients’ privacy. Considering the heavy communication overhead of Paillier, the work in [34] proposed a privacy-preserving FL using CKKS (i.e., Cheon-Kim-Kim-Song) that reduces computational overhead associated with ciphertexts. This is because CKKS is more efficient and better suitable to handle large-scale vector and multi-parameter network models compared to Paillier [35]. However, the aforementioned schemes overlook the threat of data poisoning attacks caused by distributed training.
Resisting data poisoning attacks in FL: Tolpegin et al. demonstrated that data poisoning attacks can severely reduce the classification accuracy of FL, even with a small percentage of malicious clients [36]. Additionally, they proposed a defense strategy during aggregation that identifies malicious clients to circumvent data poisoning attacks. Similarly, the works in [37, 38, 39] introduced detection mechanisms to filter malicious model updates. Differently, Doku et al. employed Support Vector Machine (SVM) to audit client’s local training data for excluding malicious clients [40]. However, this work violates the privacy of clients to some extent. Furthermore, the threats of privacy and data poisoning attacks usually coexist in practice across-silo FL. While PPFL approaches employ cryptographic primitives to ensure client privacy, they mask data poisoning attacks from malicious clients.
Resisting data poisoning attacks in PPFL: Considering the threats of both privacy and data poisoning attacks, the works such as [15, 14, 16, 17, 18] integrate HE and SMC to detect anomalous model updates directly within ciphertexts. For instance, the work in [15] proposed a privacy-enhanced FL framework that adopts HE as the underlying technology and provides two servers to punish malicious clients via gradient extraction of logarithmic function. Similarly, the work in [14] designed a validity checking protocol for ciphertexts under two servers, which protects data privacy and adjust the weight of clients’ gradients to weaken data poisoning attacks. Notably, these works [15, 14, 16, 17, 18] adopt a two-server architecture instead of a single server, which facilitates secure aggregation in adversarial scenarios. However, they neglect the deviations caused by Non-IID data, making it difficult to distinguish whether the deviations come from malicious manipulation or Non-IID data.
To audit malicious updates in PPFL with Non-IID data, the works in [19][20] are proposed to eliminate the deviations caused by Non-IID data during the aggregation stage. Specifically, ShieldFL designed a Byzantine-tolerant aggregation mechanism to prevent misjudgments on outliers caused by Non-IID data [19]. Furthermore, Chen et al. adopt clustering combined with cosine similarity and median strategies to eliminate deviations among model updates during aggregation auditing [20]. These schemes can only resist model poisoning attacks confronted by federated learning with Non-IID data, but they cannot essentially improve the performance of FL on Non-IID data. They need to be combined with specialized training techniques designed for Non-IID data (e.g., FedProx[41], FedDyn[42], or FedLC[43], etc.) to radically improve performance of FL. These specialized techniques introduce auxiliary terms into local optimization function to constrain client model updates, thereby improving consistency of model updates during local model training. However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, generating poisoned Non-IID data. This manipulation misleads specialized training techniques to adjust model updates based on these compromised inputs, which ultimately cause the global model to converge to a malicious objective. Therefore, the above works [19][20] cannot effectively integrate these specialized techniques to resist data poisoning attacks while enhancing the FL performance under poisoned Non-IID data.
Federated prototype learning: Recently, prototype learning has been applied in federated learning to address the Non-IID issue. Specifically, the works in [23, 25, 24, 26] are one of the first to propose federated prototype learning using the concept of prototype learning. Different from the specialized training techniques (e.g., FedProx, FedDyn, and FedLC), the core idea enables clients to pull the same-class samples towards the global prototypes of that class and away from the global prototypes of other classes. In other words, each class holds its corresponding prototype that is independent of other classes. Consequently, client-submitted prototype is affected by the samples but is independent of data distribution among clients. This insight motivates our work.
Although previous works [23, 25, 24, 26] have employed prototype learning to enhance the FL performance on Non-IID data, they are vulnerable to data poisoning attacks in distributed scenarios. Specifically, malicious clients can compromise the performance of federated training by uploading their poisoned prototypes. Furthermore, existing work [44] designed a dynamic memory model inversion attack that can recover the private training data by utilizing client’s learned prototypes, which seriously damages the client’s privacy. To this end, our proposed PPFPL mitigates the performance degradation caused by data poisoning attacks on Non-IID data while ensuring client privacy.
3 PRELIMINARIES
This section introduces prototypes in federated learning and CKKS technology. In addition, TABLE 1 provides explanations of the acronyms used in this paper.
Table 1: Summary of main acronyms.
Notation Meaning Notation Meaning
FL Federated Learning PPFL Privacy-Preserving Federated Learning
HE Homomorphic Encryption CKKS Cheon–Kim–Kim–Song encryption
SMC Secure Multi-Party Computation DP Differential Privacy
KGC Key Generation Center SVM Support Vector Machine
Non-IID
Non-Independent and
Identically Distributed
RSA
Rivest–Shamir–Adleman cryptosystem
3.1 Prototypes Meet Federated Learning
In the classification task of prototype learning, the prototype is defined as a feature vector representing a specific class [45]. This inherent property ensures that prototypes of the same class are similar in FL task. Consequently, many FL schemes [23, 25, 24, 26] employ prototype learning to address Non-IID challenges, enabling clients to align their local prototypes with global class representations during local model training.
To understand the prototype calculation in federated prototype learning, we introduce some basic notations below. Let $\mathcal{S}$ be the set of clients, where each client $m\in\mathcal{S}$ owns a private dataset, denoted as $\mathcal{D}_{m}=\{(\textit{{x}}_{(i)},y_{(i)})\}^{|\mathcal{D}_{m}|}$ . Here, $|\mathcal{D}_{m}|$ is the number of samples in client $m$, and $(\textit{{x}}_{(i)},y_{(i)})$ denotes sample $i$ in dataset, where $\textit{{x}}_{(i)}$ and $y_{(i)}$ correspond to the feature vector and class label of sample $i$, respectively. Meanwhile, let $\mathcal{I}$ be the set of classes in classification task, where each class $k$ belongs to $\mathcal{I}$. In classification task, the local model includes a feature extractor and a decision classifier. Specifically, the feature extractor transforms raw sample features into compressed representations, while the decision classifier maps the compressed features to get classification results. Formally, let $f_{m}(\textit{{r}}_{m,t};\cdot)$ be feature extractor for client $m$, parameterized by $\textit{{r}}_{m,t}$, where $t$ denotes the $t$-th communication round. Given the feature $\textit{{x}}_{(i)}$ of sample $i$, it is input into feature extractor to obtain compressed feature $\textit{{u}}_{(i)}=f_{m}(\textit{{r}}_{m,t};\textit{{x}}_{(i)})$. Let $g_{m}(\textit{{z}}_{m,t};\cdot)$ be decision classifier for client $m$, parameterized by $\textit{{z}}_{m,t}$. The classifier maps the compressed feature $\textit{{u}}_{(i)}$ to predict the class $y^{\prime}=g_{m}(\textit{{z}}_{m,t};\textit{{u}}_{(i)})$. Thus, we denote the local model as $\mathcal{F}_{m}((\textit{{r}}_{m,t},\textit{{z}}_{m,t});\cdot)=g_{m}(\textit{{z}}_{m,t};\cdot)\circ f_{m}(\textit{{r}}_{m,t};\cdot)$, where $\circ$ denotes composition operator. For simplicity, we use $\textit{{w}}_{m,t}$ to denote $(\textit{{r}}_{m,t},\textit{{z}}_{m,t})$, so we have $\mathcal{F}_{m}((\textit{{r}}_{m,t},\textit{{z}}_{m,t});\cdot)=\mathcal{F}_{m}(\textit{{w}}_{m,t};\cdot)$, and $\textit{{w}}_{m,t}$ is considered as model parameters for client $m$. Next, we present calculation process of prototypes.
In federated prototype learning, prototypes can be categorized into local prototypes (computed by clients) and global prototypes (aggregated by the server). Specifically, each client’s goal is to align its local prototypes with global prototypes during local model training. Each client computes its local prototype via its training dataset during local model training. Formally, let $\textit{{c}}^{k}_{m,t}$ be the local prototype of class $k\in\mathcal{I}$ at client $m$ in $t$-th communication round, calculated as:
$\begin{aligned} \textit{{c}}^{k}_{m,t}=\frac{1}{|\mathcal{D}^{k}_{m}|}\sum_{(\textit{{x}}_{(i)},y_{(i)})\in\mathcal{D}^{k}_{m}}f_{m}(\textit{{r}}_{m,t};\textit{{x}}_{(i)}),\forall k\in\mathcal{I},\end{aligned}$
where $\mathcal{D}^{k}_{m}$ denotes the dataset with class $k$ at client $m$. $\textit{{c}}^{k}_{m,t}$ can be understood as the mean of compressed features of samples belonging to class $k$ at client $m$. Further, the set of local prototypes for all classes at client $m$ is denoted as $\{\textit{{c}}^{k}_{m,t}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ . After completing local training, each client submits its local prototypes to the server for aggregation.
To calculate global prototype, the server adopts an averaging operation on local prototypes submitted by clients. Thus, the global prototype $\textit{{C}}^{k}_{t+1}$ for each class is calculated as follows:
$\begin{aligned} \textit{{C}}^{k}_{t+1}&=\frac{1}{|\mathcal{S}|}\sum_{m\in\mathcal{S}}\textit{{c}}_{m,t}^{k},\forall k\in\mathcal{I},\end{aligned}$
where $|\mathcal{S}|$ denotes the number of clients. The set of global prototypes for all classes is denoted as $\{\textit{{C}}^{k}_{t+1}\}_{k\in\mathcal{I}}^{|\mathcal{I}|}$ . Subsequently, the server distributes the latest global prototypes to each client to further train their local model.
3.2 CKKS Technique
A ciphertext is the encrypted form of plaintext data, generated through an encryption algorithm. Homomorphic encryption allows operations to be performed directly on ciphertexts without decryption. However, traditional HE schemes such as RSA [46], ElGamal [47], and Paillier are limited in that they only support either additive or multiplicative operations but not both simultaneously. In contrast, the CKKS technology provides both additive and multiplicative homomorphic encryption, which is known as full HE [48]. In addition, CKKS is known for its efficiency, especially in terms of encryption/decryption speed when applied to large-scale vectors with varying parameter lengths. Therefore, we employ CKKS to ensure client privacy while maintaining computational efficiency. The CKKS scheme mainly consists of key generation, encoding, encryption, addition, multiplication, decryption, and decoding. Here is a brief description of each operation:
  1. 1.
    Key generation. $\textit{KeyGen}(\cdot)$: Given a security parameter $\kappa$, the $\textit{KeyGen}(\cdot)$ generates secret key $Sk$, public key $Pk$ and evaluation key for ciphertext calculations.
  2. 2.
    Encoding. $\textit{Ecd}(\cdot)$: Given a $(\frac{N}{2})$-dimensional vector t and a scaling factor, the $\textit{Ecd}(\cdot)$ encodes the factor t into a polynomial $\mu$.
  3. 3.
    Encryption. $\textit{Enc}(\cdot)$: Given a polynomial $\mu$, the $\textit{Enc}(\cdot)$ encrypts $\mu$ by the public key $Pk$ and generates a ciphertext $\psi$.
  4. 4.
    Addition. $\textit{Add}(\cdot)$: Given a pair of ciphertexts $\psi_{1}$ and $\psi_{2}$, the $\textit{Add}(\cdot)$ outputs a ciphertext $\hat{\psi}=\psi_{1}\oplus\psi_{2}$, where $\hat{\psi}$ is the ciphertext of the sum of plaintexts of $\psi_{1}$ and $\psi_{2}$.
  5. 5.
@@ 3.2 CKKS Technique
  6. 6.
    Decryption. $\textit{Dec}(\cdot)$: Given a ciphertext $\psi$, the $\textit{Dec}(\cdot)$ generates a polynomial $\mu$ by the secret key $Sk$.
  7. 7.
    Decoding. $\textit{Dcd}(\cdot)$ : Given an input polynomial $\mu$ and a scaling factor $s$, the $\textit{Dcd}(\cdot)$ outputs the vector t.
For more details on the implementation principles and procedures of CKKS, please refer to the work in [49].
4 Problem Statement